[Hallc_running] [New Logentry] Hall C RC update: 10/16/2021

gaskelld at jlab.org gaskelld at jlab.org
Sat Oct 16 08:25:01 EDT 2021

Logentry Text:
The last 24 hours have been less productive than usual due to a few issues on swing and owl shift.  A water leak affected 2L02 starting at about 5 pm yesterday - we were back to production a little after 8 pm.  Unfortunately, at about 9 pm a valve in the north linac indicated it was closed (VBV1l07B) and would not indicate open after cycling.  A couple other issues came up during the downtime for the valve (RF separator trip and chopper issues), but both were resolved before completion of the valve troubleshooting, so there was no additional down time.  Hall C was back to production by about 5:30 am this morning.

We also saw a couple occurrences of the rapid LHe level drop in the bender: once on swing (https://logbooks.jlab.org/entry/3925569) and once on owl (https://logbooks.jlab.org/entry/3925713).  Heidi made some adjustments to the cryo controls to try and compensate until we can take an opportunistic access either Monday or Wednesday.  If the incidents continue, though, an access may be required sooner to troubleshoot.

The runplan continues to be production on LH2 at 42 uA with intermittent dummy running.

Reminder: there will be no RC meeting this morning. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3925852
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211016/7b285e4e/attachment.html>

More information about the Hallc_running mailing list